Write The rational number that does not have a reciprocal.

Write 
The rational number that does not have a reciprocal. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The rational number that does not have a reciprocal is 0.
Reason:
0=0/1Reciprocal of 0 = 1/0, which is not defined.

Thanks